Please help me to solve this exercise

EXERCISE 2 Parbleu Corporation, which operates in France at full capacity, produces two different p and W. In its cost accoun basis of direct labor-hours worked at full capacity, produces two different products: ting system manufacturing overhead costs are applied to products on the Basics information relating to the companys two productsis shown below ProductVV 4,500 $200 Units sold: Average selling price: Units produced: Direct materials (per unit): Direct labors (per unit): Machine hours (per unit) Sales commission ProductZ” 8,000 $250 7,000 $60.00 5 hours 3 hours 5% 5,000 $50.00 2,5 hours 6 hours 4% The company has provided the following financial data concerning its actual operations for the year ended: Direct materials purchased Direct labor Manufacturing Fixed Overhead Manufacturing Variable Overhead Sales General& Administrative Costs (fixed) Interest Revenues $ 695,000 $ 593,750 $608.000 S 35,625 530,000 $24,700 The initial inventory for Direct material at the beginning of the year was $ 80,000. Final inventory is $105,000. You are required: a) b) to determine the cost of the two products per unit using absorption costing to provide a segmented income statement in which contribution margin and EBIT for the company are shown
Answer #1
a) Product Z Product W Total
Direct materials (7000*60+5000*50) 420000 250000 670000
Direct labor (7000*5*12.5/5000*2.5*12.5) 437500 156250 593750
Variable overhead (7000*5*0.75/5000*2.5*0.75) 26250 9375 35625
Fixed overhead (7000*5*12.8/5000*2.5*12.8) 448000 160000 608000
Total cost of production under absorption costing 1331750 575625 1907375
Number of units produced 7000 5000
Cost per unit $        190.25 $        115.13
Working Notes:
Total direct labor hours = 7000*5+5000*2.5 = 47500
Direct labor hour rate = 593750/47500 = $          12.50
Direct materials used = 80000+695000-105000 = 670000
Variable overhead allocation rate = 35625/47500 = $            0.75 per DLH
Fixed overhead allocation rate = 608000/47500 = $          12.80 per DLH
b) Product Z Product W Total
Sales revenue (8000*250/4500*200) 2000000 900000 2900000
Variable cost of goods sold 883750 415625 1299375
Contribution margin 1116250 484375 1600625
Fixed manufacturing costs 608000
Sales general and administrative costs (fixed) 530000
EBIT ` 462625
Interest revenues 24700
EBT 487325